OpenStudy (anonymous):

Solve for y: (please show steps) 6x+3y=9

OpenStudy (lgbasallote):

first...start by subtracting both sides by 6x. what do you get?

OpenStudy (anonymous):

X+3Y=3

OpenStudy (lgbasallote):

hmm what did you do?

OpenStudy (anonymous):

ohh oops! i forgot to the x part. i misread your directions

OpenStudy (anonymous):

3y=9-6x

OpenStudy (lgbasallote):

ahh good. now divide everything by 3

OpenStudy (anonymous):

y=3-2x

OpenStudy (lgbasallote):

that's the answer now

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome